What is the time period of object to displace from mean position to half displacement position?

Grade:12th Pass

1 Answers

areeba
6 years ago
As x =Xm/2 so using the equation x =Xm sinθt .putting the value of x then Xm/2=Xm sinθt .simplyfy the. Equation 1/2=sinθt so (1/2 )sin inverse =θt. next pi/6 or 30 degree =θt where θ=wt and w=2pi/T so pi/6 =2pi/T (t) simplifying the equatin we obtain the ans t =T/12